local time is 3 am and the longitude is 120 east what time is the prime meridian


1
Expert's answer
2021-02-03T09:06:20-0500

1 hour difference corresponds to 15° difference in longitude. Thus, 120° correspond to


"t=\frac{120}{15}=8\text{ h}."

The longitude is 120° to the east from the Prime meridian, which means that 3 am at 120° is 8 h ahead of the time at the Prime meridian:


"t_\text{P.M.}=24-(8-3)=19\text{ h},"

or 7 pm.